OpenStudy (anonymous):

The distance across the circular mouth of a volcano running through its center is 75 yrds. What is the distance around its mouth?

OpenStudy (anonymous):

75/2 = 27.5 circumference = 2*r*pi = 75 * pi = 235.62

OpenStudy (anonymous):

thanks